# Assessing the accuracy of the star formation rate measurements by direct star count in molecular clouds

### v2

@@ 5 Conclusions
In this work, we explored under which circumstances counting YSOs in star-forming regions can yield reliable estimates of the SFR. To this end, we developed a Monte Carlo models in which the masses, the birth times of the YSOs along with the lifetimes associated with the different YSO Classes (Class 0, I, F, II, and III) are all randomly drawn from distributions function that describe each of these quantities. The masses of the protostars are randomly drawn from an IMF which can be either similar to that of the Milky Way field or be variable in the range of variations observed among young clusters in the Galaxy. The birth times of protostars are either drawn from a flat distribution function (i.e., constant SFH), or from a time-dependent, burst-like function. Finally, the lifetime of the YSOs associated with each protostellar class are randomly drawn from Gaussian distribution functions centered around those suggested by the observations (Evans et al. 2009). Using these prescriptions, we follow the time evolution of the number of YSOs in different evolutionary classes and calculate the corresponding SFR at every epoch. We find that the SFRs derived using the Class 0 population reproduce the true SFR, and this conclusion is valid irrespective of the shape of the SFH. For a constant SFH, the SFR derived using the more evolved populations of YSOs (Class I, Class F, Class II, and Class III) reproduce the real SFR at later epochs. For example for an SFR estimate that is based on the Class II population and lifetimes, the real SFR is reproduced after Myr from the beginning of star formation. For a time-dependent burst-like SFH, all SFR estimates based on the number counts of the more evolved populations (all classes but Class 0) fail to reproduce the true SFR. Also, we find that SFR estimates that are based on the total number of YSOs associated with a characteristic lifetime for all populations fail to reproduce the real SFR, irrespective of the shape of the SFH. Furthermore, we show that all of these conclusions are independent of the shape of stellar initial mass function.
@@
The synthetic models presented in this work can help shed light on the SFH of observed star-forming regions. We show that the models with a constant SFH make different predictions for the SFR offsets vs. the number ratios of YSOs than those of a burst-like SFH. For five, nearby, low-mass star-forming regions that constitute the Evans et al. 2009 sample (i.e, the Cha II, Lupus, Perseus, Serpens, and Ophiucus clouds), we show that the relations between the SFR offset parameters and of the clouds with the number ratios and , respectively, are matched by models with a constant SFR. A comparison between our models and a larger number of star-formation regions, such as those that would be observed by the ABYSS project (Kounkel et al. 2023) and with the James Webb Science Telescope (e.g., Lenkić et al. 2024; Peltonen et al. 2024) will enable us to distinguish between the SFHs of star forming clouds as a function of some of their fundamental properties such mass and surface density. In this work, we have not accounted for the effects of mass accretion which induce variability in their luminosity and may lead to misclassification between YSO classes. We believe however that such effects will not change our conclusion as only a small fraction of the YSO population shows sign of variability (Mairs et al. 2017).
@@
Deriving accurate SFRs is crucial for the interpretation of the scaling laws of star formation on individual molecular cloud scales and for testing both theories and numerical simulations star formation (e.g., Hennebelle & Chabrier 2011; Dib 2011; Padoan & Nordlund 2011; Federrath & Klessen 2012; Burkhart 2018; Eden & Teyssier 2024). In particular, the SFR-surface gas density () relation on local scales is observed to be steeper than the one for entire galaxies or the one derived on kpc scales within galaxies (e.g., Heiderman et al. 2010; Hony et al. 2015; Pokhrel et al. 2020,2021). One interpretation of this steep relation between and the SFR is that the higher SFR observed in individual regions is due to the selection bias introduced by studying a region which is actively forming stars, while the galaxy-averaged values include regions which are forming stars at different rates (Kruijssen & Longmore 2014). Another interpretation is that the steep -SFR relation for individual clouds is the reflection of an evolutionary sequence in which the SFRs of contracting clouds increase much faster than their mean densities or mean surface densities and this leads to a steep supra-linear -SFR relation (Zamora-Avilés et al. 2012). Further work is needed in order to disentangle these effects from the offsets in the SFR estimates that we have uncovered in this work and which are solely due to biases introduced by the star counting method.

### v4

@@ 5 Conclusions
In this work, we explore the circumstances under which counting YSOs in star-forming regions can yield reliable estimates of the SFR. To this end, we developed a Monte Carlo model in which the masses, the birth times of the YSOs, and the lifetimes associated with the different YSO classes (Classes 0, I, F, II, and III) are all randomly drawn from distribution functions that describe each of these quantities. The masses of the protostars are randomly drawn from an IMF that can be either similar to that of the Milky Way field or variable in the range of variations observed among young clusters in the Galaxy. The birth times of protostars are either drawn from a flat distribution function (i.e., constant SFH), or from a time-dependent, burst-like function. Finally, the lifetimes of the YSOs associated with each protostellar class are randomly drawn from Gaussian distribution functions centered around those suggested by the observations (Evans et al. 2009). Using these prescriptions, we follow the time evolution of the number of YSOs in different evolutionary classes and calculate the corresponding SFR at every epoch. We find that the SFRs derived using the Class 0 population reproduce the true SFR, and this conclusion is valid irrespective of the shape of the SFH. For a constant SFH, the SFRs derived using the more evolved populations of YSOs (Class I, Class F, Class II, and Class III) reproduce the real SFR at later epochs. For example, for an SFR estimate that is based on the Class II population and lifetimes, the real SFR is reproduced after Myr from the beginning of star formation. For a time-dependent burst-like SFH, all SFR estimates based on the number counts of the more evolved populations (all classes but Class 0) fail to reproduce the true SFR. Also, we find that SFR estimates that are based on the total number of YSOs associated with a characteristic lifetime for all populations fail to reproduce the real SFR, irrespective of the shape of the SFH. Furthermore, we show that all of these conclusions are independent of the shape of the stellar IMF.
@@
The synthetic models presented in this work can help shed light on the SFH of observed star-forming regions. We show that the models with a constant SFH make different predictions for the SFR offsets versus the number ratios of YSOs from those with a burst-like SFH. For five, nearby, low-mass star-forming regions that constitute the Evans et al. 2009 sample (i.e, the Cha II, Lupus, Perseus, Serpens, and Ophiucus clouds), we show that the relations between the SFR offset parameters and of the clouds and the number ratios and , respectively, are matched by models with a constant SFR. A comparison between our models and a larger number of star-formation regions, such as those that would be observed by the ABYSS project (Kounkel et al. 2023) and with the James Webb Science Telescope (e.g., Lenkić et al. 2024; Peltonen et al. 2024) will enable us to distinguish between the SFHs of star-forming clouds as a function of some of their fundamental properties, such as mass and surface density. In this work, we have not accounted for the effects of mass accretion, which induces variability in the luminosity of YSOs and may lead to their misclassification. We believe however that such effects will not change our conclusion, as only a small fraction of the YSO population shows signs of variability (Mairs et al. 2017).
@@
Deriving accurate SFRs is crucial for the interpretation of the scaling laws of star formation on individual molecular cloud scales and for testing both theories and numerical simulations of star formation (e.g., Hennebelle & Chabrier 2011; Dib 2011; Padoan & Nordlund 2011; Federrath & Klessen 2012; Burkhart 2018; Eden & Teyssier 2024). In particular, the SFR–surface gas density () relation on local scales is observed to be steeper than the one for entire galaxies or the one derived on approximately kiloparsec scales within galaxies (e.g., Heiderman et al. 2010; Hony et al. 2015; Pokhrel et al. 2020,2021). One interpretation of this steep relation between and the SFR is that the higher SFR observed in individual regions is due to the selection bias introduced by studying a region that is actively forming stars, while the galaxy-averaged values include regions that are forming stars at different rates (Kruijssen & Longmore 2014). Another interpretation is that the steep –SFR relation for individual clouds is the reflection of an evolutionary sequence in which the SFRs of contracting clouds increase much faster than their mean densities or mean surface densities and this leads to a steep supra-linear –SFR relation (Zamora-Avilés et al. 2012). Further work is needed in order to disentangle these effects from the offsets in the SFR estimates uncovered in the present work, which are solely due to biases introduced by the star counting method.
@@ Acknowledgements.
- Jose (2013) Jose, J., Pandey, A. K., Samal, M. R. et al. 2013, MNRAS, 432, 3445
- Jose (2016) Jose, J., Kim, J. S., Herczeg, G. J. et al. 2016, 822, 49